OpenStudy (anonymous):

The displacement (in centimeters) of a particle moving back and forth along a straight line is given by the equation of motion s = 3 sin πt + 5 cos πt, where t is measured in seconds. (Round your answers to two decimal places.) (a) Find the average velocity during each time period. [1,2] [1,1.1] [1,1.01] [1,1.001] (b) Estimate the instantaneous velocity of the particle when t = 1.

OpenStudy (adamaero):

Differentiate.

OpenStudy (anonymous):

i'm sorry what do you mean?

OpenStudy (adamaero):

ds / dt similar to dy/dx ~does that look familiar?

OpenStudy (anonymous):

not at all... if you could explain it and help me work through the problems that would be helpful.

OpenStudy (adamaero):

what math are you in

OpenStudy (anonymous):

calc

OpenStudy (adamaero):

Ok, what is the title of the chapter you're on?

OpenStudy (adamaero):

Differentiation, so you should read it

OpenStudy (anonymous):

2.1 is the chapter

OpenStudy (adamaero):

also you should memorize trig identities (d/dx)(sin(x)) = cos(x) (d/dx)(cos(x)) = -sin(x)

OpenStudy (anonymous):

okay I'm opening my textbook up now to take another look

OpenStudy (adamaero):

d/dx of position (displacement) = velocity d/dx of velocity = acceleration d/dx of acceleration = jerk etc... snap crackle pop jouce

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I'm looking at my textbook and it still is making no sense

OpenStudy (adamaero):

lol

OpenStudy (adamaero):

this guy is the best: https://www.youtube.com/watch?v=pFeuGMMiZWw

OpenStudy (adamaero):

she's also pretty good: https://www.youtube.com/watch?v=4GaxepQ2j-s